Consider the following random sample: 16, 13, 10, 17, 16, 14, 5, 12, 15, 8. Calculate
the first quartile (choose the best answer).
A. 9.5 B. 7.5 C. 10.5 D. 0.25 E. 11
Answer #1

Solution:
Arranging Observations in the ascending order, We get :
5,8,10,12,13,14,15,16,16,17

Here, n=10

Q1=(n+14)th value of the observation

=(11/4)th value of the observation

=(2.75)th value of the observation

=2nd observation +0.75[3rd-2nd]

=8+0.75[10-8]

=8+0.75(2)

=8+1.5

=9.5

Q1= 9.5

option A ) is correct.